# Soft-TransFormers for Continual Learning

### v2

@@ 4.1 Soft-MSA Layers
To address the task-specific fine-tuning of the pre-trained model, such as ViT, this work proposes a new Soft-TransFormer (Soft-TF), as illustrated in Figure 1. The proposed Soft-TF consists of a conventional neural network, like a multilayer transformer with multihead attention and forward networks. Using well-trained transformer parameters, we could discover task-specific soft-networks, as depicted in Figure 4. The Soft-TF incrementally learns model weights and task-adaptive soft-masks with well-pre-trained and soft-network parameters $\bm{m}$.
The Soft-TF originates from learned parameters distributed with $\mu\approx 1.0$ & various variances, as stated in Figure 3: the histogram density plots compare the distribution of parameters for the attention layer L[12] in two contexts: the masking mechanism ($\bm{m}^{QKV}$) and the fine-tuned weighted product ($\bm{w}^{QKV}\odot\bm{m}^{QKV}$) for two tasks (Task 2 and Task 3). In Figure 3(a), the $\bm{m}^{QKV}$ parameters are concentrated in a narrow range centered around 1. There is a slight variation between Task 2 and Task 3. However, the overall range remains consistent. $\bm{m}^{QKV}$ acts as a soft gating mechanism, controlling which pre-trained parameters are retained or modified during fine-tuning. In LABEL:{fig:soft_weight_cifar10}(b), the distribution of pre-trained weight ($\bm{w}^{QKV}$) is centered around 0, with a Gaussian-like spread. The distribution of soft-masked fine-tuned weight ($\bm{w}^{QKV}\odot\bm{m}^{QKV})$ remain similar to that of $\bm{w}^{QKV}$ but slightly narrower. The overlap between the two distributions indicates that the fine-tuning process maintains the general structure of the pre-trained weights while introducing task-specific modifications. By using soft-masks $\bm{m}^{QKV}$, Soft-TF ensures task-specific fine-tuning, reducing the risk of catastrophic forgetting and improving performance in various sequential tasks. The task-specific Soft-TF fine-tuning enables the models to infer task ID ideally in following Vision CIL scenarios.
Refer to caption Refer to caption
(a) $\bm{m}^{QKV}$ (b) $\bm{w}^{QKV}$ v.s. $\bm{w}^{QKV}\odot\bm{m}^{QKV}$
Given a pre-trained parameter $\bm{\theta}$ and learnable soft-parameters $\bm{m}$, Soft-ViT is represented as $f_{\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}$, consisting of $N$ consecutive soft-MSA layers. We extend the notation by denoting the input embedding feature of the $l_{\ast}$-th learnable soft-MSA layer as $\bm{h}^{(l_{\ast})}$, where $l_{\ast}=1,2,\dots,N$, and $l_{\ast}$ can refer to either the G-Prompt layer $l_{g}$ or the E-Prompt layer $l_{e}$. Note that while the pre-trained parameters $\bm{\theta}$ remain fixed, the soft-parameters $\bm{m}$ are updated to provide task-specific solutions.
G-prompt. $\bm{g}\in\mathbb{R}^{L_{g}\times D}$ with sequence length $L_{g}$ and embedding dimension $D$, is a shared parameter for all tasks. G-Prompt is attached to the $l_{g}$-th MSA layer to transform $\bm{h}^{(l_{g})}$ via a prompting function as follows:
$\bm{h}_{g}^{(l_{g})}=f^{prompt}_{\bm{\theta}}(\bm{g},\bm{h}^{(l_{g})}),$
where $f^{prompt}_{\bm{\theta}}$ defines the approach for attaching the prompt to the hidden embeddings.
E-prompt & Soft-networks. $\bm{e}=\{\bm{e}_{t}\}^{\mathcal{T}}_{t=1}$ is a set of task-dependent parameters, where $\bm{e}_{t}\in\mathbb{R}^{L_{e}\times D}$ has as sequence length of $L_{e}$ and the same embedding dimension $D$ as the G-prompt, and $\mathcal{T}$ is the total number of tasks. Unlike the shared G-prompt, each $\bm{e}_{t}$ is associated with a task-specific key $\bm{k}_{t}\in\mathbb{R}^{D}$, which is also a learnable parameter aimed at capturing representative features of a task. For an input example from the $t$-th task, to attach E-prompt to the $l_{e}$-th soft-MSA layer, we apply the prompting function in a similar way:
$\bm{h}_{e}^{(l_{e})}=f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}(\bm{e}_{t},\bm{h}^{(l_{e})}).$
Figure 4: Soft-TransFormers (Soft-TF): At training time, the E-Prompt and the Soft-network are selected according to task identity, and the selected G-Prompt, E-Prompt, and the Soft-networks (Soft-Attention and Feed Forwards) are trained together with a classifier. At test time, an input is transformed by a query function (Prompt ID) or task identifier (Gradient ID) to match the closest task key $\bm{k}_{t}$, E-prompt $\bm{e}_{t}$ and Soft-networks $\bm{m}_{t}^{\{K,Q,V\}}$. Note task identifier is depicted in Section 5.
4.2 Prompts with Learnable Subnetworks
G- and E-prompts, along with learnable soft-networks, encode specific types of instructions during training with the backbone and work together to guide the model’s predictions during inference. We have demonstrated the method for attaching prompts and learnable soft-networks to a single soft-MSA layer. Similarly to the approach taken in DualPrompt (Wang et al., 2022b), we also investigate layers of E-prompts with learnable soft-networks $\bm{m}$, while utilizing the layers designated for G-prompts.
Layers of G- and E-Prompts. We use the multilayered extension of both types of prompts: $\bm{g}=\{\bm{g}^{(l_{g})}\}_{l_{g}=start_{g}}^{end_{g}}$, where $\bm{g}^{(l_{g})}\in\mathbb{R}^{L_{g}\times D}$ represents the G-prompt attached to the $l_{g}$-th MSA layer. Similarly, we define $\bm{e}_{t}=\{\bm{e}_{t}^{(l_{e})}\}_{l_{e}=start_{e}}^{end_{e}}$ for the $l_{e}$-th conventional MSA layer. In this configuration, the G-prompt $\bm{g}^{(l_{g})}$ is attached from the $start_{g}$-th to the $end_{g}$-th conventional MSA layers, and the E-prompt $\bm{e}_{t}^{(l_{e})}$ is attached to the $\left[start_{e},end_{e}\right]$-th soft-MSA layers, ensuring that there is no overlap between them. In our experiments, we follow the $l_{g}\notin\left[start_{e},end_{e}\right]$ ($l_{g}=[1,2]$) settings used in DualPrompt and empirically search for the optimal $\left[start_{e},end_{e}\right]$ layers for the learnable subnetworks through ablation studies.
Learnable Soft-networks. The prompting function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ determines how prompts ($\bm{p}$) are combined with fine-tuned soft ($\bm{\theta}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}$) embedding features. From another perspective, $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ directly influences the interaction between high-level instructions in the prompts and low-level representations. Therefore, we believe that a well-designed prompting function, along with task-specific parameters, is crucial for optimizing overall continual learning performance.
Specifically, applying a prompting and fine-tuning function $f^{prompt}_{\bm{\theta\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}}}}$ can be seen as modifying the inputs to the soft-MSA layers. Let the input to the soft-MSA layer be $\bm{h}\in\mathbb{R}^{L\times D}$, and denote the input query, key, and values for the soft-MSA layer as $\bm{h}_{Q}$, $\bm{h}_{K}$, and $\bm{h}_{V}$, respectively. A soft-MSA layer is defined by the following equation:
$\text{MSA}(\bm{h}_{Q},\bm{h}_{K},\bm{h}_{V})=[\bm{h}_{1},\cdots,\bm{h}_{i},\cdots,\bm{h}_{n}]\bm{w}^{O}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}}$
where $\bm{h}_{i}=\text{Attention}(\bm{h}_{Q}(\bm{w}^{Q}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}}),\bm{h}_{K}(\bm{w}^{K}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}),\bm{h}_{V}(\bm{w}^{V}_{i}\odot{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}}))$; $\bm{w}_{i}^{O},\bm{w}_{i}^{Q},\bm{w}_{i}^{K}$, and $\bm{w}_{i}^{V}$ are fixed projection matrices while ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{O}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{Q}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{K}}$, and $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\bm{m}^{V}$} are learnable parameters. $s$ is the number of heads. In ViT, $\bm{h}_{Q}=\bm{h}_{K}=\bm{h}_{V}$. Here, we define a unified prompt parameter with a sequence length of $L_{p}$, such as $\bm{p}\in\mathbb{R}^{L_{p}\times D}$ for a single-layered G- or E-prompt.
4.3 Fine-tuning on Well-initialized Parameters
In this framework, we concatenate the prompts $\bm{p}_{t}$ and the embedding sequence $\bm{x}_{t}$, i.e., inputs from $t$-th task, along the embedding dimension: $\bm{z}_{t}=[\bm{p}_{t};\bm{x}_{t}]$. With the weights of $\bm{w}^{Q}\odot\bm{m}^{Q},\bm{w}^{K}\odot\bm{m}^{K},\bm{w}^{V}\odot\bm{m}^{V}$, the soft-transformer takes query ($\bm{q}_{t}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$) and key ($\bm{k}_{t}=(\bm{w}^{K}\odot\bm{m}^{K})\bm{z}_{t}$) as input of the soft-MSA layer. The soft-attention matrix is then given by:
$\bm{a}_{t}=\text{softmax}\left(\frac{\bm{q}_{t}\bm{k}_{t}^{T}}{\sqrt{D/n}}\right)$
where we focus on $\bm{q}_{t}\bm{k}_{t}^{T}=(\bm{w}^{Q}\odot\bm{m}^{Q})\bm{z}_{t}$ and $\bm{z}_{t}^{T}(\bm{w}^{K}\odot\bm{m}^{K})^{T}$. First, the trainable prompt parameters can be denoted as:
$\bm{z}_{t}\cdot\bm{z}_{t}^{T}=\begin{bmatrix}\bm{p}_{t}\\ \bm{x}_{t}\end{bmatrix}\begin{bmatrix}\bm{p}_{t}\;\bm{x}_{t}\end{bmatrix}=\begin{bmatrix}\bm{p}_{t}\bm{p}^{T}_{t}\;\bm{p}_{t}\bm{x}_{t}^{T}\\ \bm{x}_{t}\bm{p}_{t}^{T}\;\bm{x}_{t}\bm{x}_{t}^{t}\end{bmatrix}$
Second, the trainable soft-attention layer’s parameters with $\bm{m}^{Q}$ and $\bm{m}^{K}$ are as follows:
$\left\{\begin{matrix}\bm{m}^{Q}\cdot\bm{p}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{p}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{p}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T},\\ \bm{m}^{Q}\cdot\bm{x}_{t}\bm{x}_{t}^{T}\cdot(\bm{m}^{K})^{T}.\\ \end{matrix}\right\}$
where $\bm{w}^{Q}$ and $\bm{w}^{K}$ are frozen and unchanged during training and test. The trainable soft-attention layer’s parameters $\bm{m}^{\ast}$ are used to analyze the convergence rate of the Soft-Transformer (see Section A.1 stated in detail).
4.4 Optimization of Soft-Transformers
The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm 1 and Algorithm 2. We denote the architecture with attached prompts as $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$. The input $\bm{x}$ of the $t$-th task is transformed using $f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}$ and then passed to the classification head $f\phi$, parameterized by $\phi$, for prediction. Finally, we train the two prompts, the task keys, the soft-attention parameters, and the newly-initialized classification head in an end-to-end manner:
$\min_{\bm{g},\bm{e}_{t},\bm{m}_{t},\bm{k}_{t},\bm{\phi}}\mathcal{L}(f_{\phi}(f_{\bm{g},\bm{e}_{t},{\bm{m}_{t}}}(\bm{x})),y)+\lambda\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t}),$
Here, $\mathcal{L}$ represents the cross-entropy loss, and $\mathcal{L}_{\text{match}}(\bm{x},\bm{k}_{t})=\gamma(\bm{q}(\bm{x}),\bm{k}_{t})$ denotes the matching loss, where $\bm{q}(\bm{x})=f(\bm{x})[0]$ corresponds to the feature vector associated with the [class] token (Dosovitskiy et al., 2020; Wang et al., 2022b), and $\gamma$ is the cosine similarity. The scalar $\lambda$ serves as a balancing factor between the losses; here, we follow the same DualPrompt setting as a baseline.
@@ 4.4 Optimization of Soft-Transformers
Table 1: Performances of Vision Class Incremental Learning (CIL, ViTs) in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note that Soft-TF performed without FeedForward ($F$) networks. $\ast$ denotes our reproduced results.
ViT-B/12 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
@@
 + Adapter-L[all], r=75 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
 + LoRA-L[10,11,12], r=500 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.21 15.09K / 53
@@
 + Soft-TF-L[10,11,12] 6.93M Gradient ID 98.05 / 0.25 13.71K / 130 98.96 / 0.23 15.51K / 163 83.70 / 0.53 15.08K / 82

### v3

@@ 4.1 Soft-MSA Layers
Let $\bm{\theta}$ denote the parameters of a pre-trained Transformer, which remain fixed throughout continual learning. For each task $t$, as depicted in Figure˜3, Soft-TF introduces a set of real-valued masks
$\bm{m}_{t}=\{\bm{m}_{t}^{Q},\,\bm{m}_{t}^{K},\,\bm{m}_{t}^{V},\,\bm{m}_{t}^{O}\},$
which multiplicatively modulate the attention projections. Given hidden features $h$, we obtain:
$\bm{Q}=h(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q}),\;\bm{K}=h(\bm{\theta}^{K}\odot\bm{m}_{t}^{K}),\;\bm{V}=h(\bm{\theta}^{V}\odot\bm{m}_{t}^{V}).$
The output of the multihead attention layer is:
$\mathrm{MSA}(h)=\mathrm{Attn}(\bm{Q}_{i},\bm{K}_{i},\bm{V}_{i})\Big(\bm{\theta}^{O}\odot\bm{m}_{t}^{O}),$ (1)
where $\mathrm{Attention}(\cdot)$ denotes standard scaled dot-product attention.
Intuition. The masks are initialized at $\mathbf{1}$, ensuring that the initial computation matches the pre-trained model exactly. During optimization, the masks induce task-specific, continuous subnetwork selection within each attention projection. This allows Soft-TF to adjust representational pathways while keeping the global structure of the backbone fixed, thereby constraining parameter drift and enabling stable adaptation across tasks.
4.2 Prompts Coupled with Soft-Subnetwork
Soft-TF incorporates both general prompts (shared across tasks) and expert prompts (task-specific) into a multi-layer prompting mechanism. Let $\bm{g}=\{\bm{g}^{l}\}_{l=l_{g}^{\text{start}}}^{l_{g}^{\text{end}}}$, $\bm{e}_{t}=\{\bm{e}_{t}^{l}\}_{l=l_{e}^{\text{start}}}^{l_{e}^{\text{end}}}$ denote the layerwise G-prompts and E-prompts, respectively, with the constraint $l_{g}^{\text{start}},l_{g}^{\text{end}}]\cap[l_{e}^{\text{start}},l_{e}^{\text{end}}]=\emptyset.$
Given an input sequence $\bm{x}$, prompts are concatenated as
$\bm{z}=[\,\bm{p};\,\bm{x}\,],\quad\bm{p}\in\{\bm{g},\bm{e}_{t}\}.$
Soft-MSA then processes $\bm{z}$ via masked projections:
$\bm{q}=(\bm{\theta}^{Q}\odot\bm{m}_{t}^{Q})\bm{z},\qquad\bm{k}=(\bm{\theta}^{K}\odot\bm{m}_{t}^{K})\bm{z},$
yielding the attention matrix
$\bm{a}=\mathrm{softmax}\left(\frac{\bm{q}\bm{k}^{\top}}{\sqrt{D/s}}\right).$ (2)
G-prompts impart globally relevant cues, while E-prompts and soft-masks enable fine-grained, task-specific adjustments. This combination yields strong task separation without requiring explicit task identifiers during evaluation.
4.3 Optimization over Frozen Backbones
For task $t$, the predictive model is
$f_{t}(\bm{x})=f_{\phi}\big(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(x)\big),$
where $\phi$ is a lightweight task classifier. Training optimizes:
$\min_{\bm{g},\,\bm{e}_{t},\,\bm{m}_{t},\,\bm{k}_{t},\,\phi}\;\mathcal{L}_{\mathrm{CE}}\big(f_{\phi}(f_{\bm{\theta},\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})),y\big)+\lambda\,\gamma\big(q(\bm{x}),\bm{k}_{t}\big),$ (3)
where $\bm{k}_{t}$ is a task-level key vector and $\gamma$ is a cosine-similarity regularizer, defined as $1-\cos(\cdot,\cdot)$, which facilitates task inference at test time. Only the prompts, soft-masks, and classifier parameters are updated; the backbone remains frozen, yielding low memory cost and consistent forward-time efficiency. The overall process of the Soft-TransFormers (Soft-TF) during training and testing is described as Algorithm˜1 and Algorithm˜2.
Algorithm 1 Training Procedure of DualPrompt-Soft-TF
1:  Input: Pre-trained transformer backbone $f$, classifier $f_{\phi}$;
2:   number of tasks $\mathcal{T}$; training sets $\{\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{n_{t}}\}_{t=1}^{\mathcal{T}}$;
3:   G-Prompt $\bm{g}$, E-Prompts $\bm{E}=\{\bm{e}_{t}\}_{t=1}^{\mathcal{T}}$;
4:   task keys $\bm{K}=\{\bm{k}_{t}\}_{t=1}^{\mathcal{T}}$;
5:   soft-networks $\bm{M}=\{\bm{m}_{t}\}_{t=1}^{\mathcal{T}}$;
6:   prompt insertion layers $l^{start}_{g},l^{end}_{g}]$, $l^{start}_{e},l^{end}_{e}]$;
7:   prompting function $f_{\bm{\theta}\odot\bm{m}}$; epochs $\mathcal{K}_{t}$ for task $t$.
8:  Initialize: $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$
9:  for task $t=1,\dots,\mathcal{T}$ do
10:  Select task-specific $\bm{e}_{t}$, soft-network $\bm{m}_{t}$, and key $\bm{k}_{t}$
11:  Construct prompted backbone $f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}$ by attaching:
12:   G-Prompt to layers $l^{start}_{g},l^{end}_{g}]$,
13:   E-Prompt to soft-MSA layers $l^{start}_{e},l^{end}_{e}]$ via $f_{\bm{\theta}\odot\bm{m}_{t}}$
14:  for epoch $s=1,\dots,\mathcal{K}_{t}$ do
15:   Sample mini-batch $B=\{(\bm{x}_{i,t},y_{i,t})\}_{i=1}^{l}$
16:   for each $(\bm{x},y)$ in $B$ do
17:    Compute prompted representation:
18:     $\bm{h}=f_{\bm{g},\bm{e}_{t},\bm{m}_{t}}(\bm{x})$
19:    Compute prediction $\hat{y}=f_{\phi}(\bm{h})$
20:    Compute per-sample loss:
21:     $\mathcal{L}_{x}=\mathcal{L}_{CE}(\hat{y},y)+\lambda\,\gamma(q(\bm{x}),\bm{k}_{t})$
22:   end for
23:   Update $\phi$, $\bm{g}$, $\bm{E}$, $\bm{M}$, $\bm{K}$ via backpropagation
24:  end for
25:  end for
4.4 Stability and Bounded Parameter Drift
Define the effective parameters for task $t$ as
$\bm{w}=\bm{\theta}\odot\bm{m}_{t},$
so that $\bm{w}^{(1)}=\bm{\theta}$ under mask initialization $\bm{m}_{t}^{(1)}=\mathbf{1}$. A single gradient update yields:
$\bm{w}^{(k+1)}=\bm{w}^{(k)}-\eta\big(\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\big).$
Thus, the optimization path is governed by a rescaled gradient whose magnitude is bounded by the scale of $\bm{\theta}$. Under standard convex–Lipschitz conditions, we obtain the cumulative suboptimality bound:
$\begin{aligned} \sum_{k=1}^{T}&\langle\bm{w}^{(k)}-\bm{w}^{\star},\,\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\rangle\leq\\ &\mathcal{O}\!\left(\frac{\|\bm{w}^{(1)}-\bm{w}^{\star}\|^{2}}{\eta}+\eta\sum_{k=1}^{T}\|\bm{\theta}\odot\nabla_{\bm{m}}\mathcal{L}\|^{2}\right).\end{aligned}$
@@ 4.4 Stability and Bounded Parameter Drift
Table 1: Performances of Vision Class Incremental Learning in terms of Accuracy, Forgetting, Trainable parameters, and Time complexity on 10/20-Split-CIFAR100 and 10-Split-ImageNet-R. Note that Soft-TF performed without FeedForward ($F$) networks. $\ast$ denotes our reproduced results and L[10,11,12] represents Soft-TransFormer of 10, 11, 12 Layers.
ViT-B/16 10-Split-CIFAR100 20-Split-CIFAR100 10-Split-ImageNet-R
@@
+ Adapter 6.93M Prompt ID 86.51 / 4.75 14.98K / 90 84.48 / 5.81 16.13K / 109 70.56 / 4.71 15.65K / 54
+ LoRA 6.93M Prompt ID 82.00 / 4.33 13.24K / 79 92.14 / 2.02 15.89K / 105 43.51 / 13.2 15.09K / 53
@@
NoRGa 0.03M Prompt ID 94.48 / 1.44 12.12K / 76 97.39 / 0.58 11.60K / 78 75.40 / 4.59 13.10K / 47
+ Soft-TF-L[10,11,12] (SOTA) 6.93M Prompt ID 98.25 / 0.31 13.87K / 80 98.30 / 0.30 15.60K / 104 91.94 / 0.34 15.35K / 52
Joint of DualPrompt 0.13M - 90.85 - 90.85 - 79.13 -
Joint of DualPrompt+Soft-TF 6.93M - 95.90 - 95.90 - 80.21 -
Joint of NoRGa+Soft-TF 6.93M - 98.20 - 98.20 - 90.73 -
Implications for continual learning. The initial distance $\|\bm{w}^{(1)}-\bm{w}^{\ast}\|$ is naturally small across most tasks because the pre-trained backbone already provides strong feature representations. Consequently, by standard convergence guarantees of first-order optimization for Lipschitz-continuous objectives without assuming strong convexity, the suboptimality decreases at a rate of $\mathcal{O}(1/\sqrt{T})$. More importantly, since optimization occurs only in the low-dimensional soft-parameter space $(\bm{m}_{t}^{Q},\bm{m}_{t}^{K},\bm{m}_{t}^{V})$ while $\bm{\theta}$ remains fixed, the entire trajectory $\bm{w}^{(k)}$} stays confined to a narrow region around the pretrained solution manifold. This imposes an inherent bound on parameter drift and explains why Soft-TF reliably preserves global knowledge encoded in $\bm{\theta}$ while still enabling targeted adaptation to each new task. Taken together, these theoretical results (see Section˜A.1) provide a principled foundation for our design choice: initializing masks at $\mathbf{1}$ and fine-tuning only $\bm{m}_{t}$ over a frozen backbone yields stable optimization, bounded deviation from pre-trained representations, and substantially reduced catastrophic forgetting in continual learning.